OpenStudy (livyloo52):

Problem Page Boris is on his way home in his car. He has driven 30 miles so far, which is five-sixths of the way home. What is the total length of his drive?

OpenStudy (brainzonly):

scratch that. I think ive got an idea

OpenStudy (brainzonly):

\[\frac{ 30 }{ 1 }*\frac{ 5 }{ 6 }\]

OpenStudy (kevin):

@brainzonly can help you. He is smart :D

OpenStudy (brainzonly):

wow, abandon me. :( 30*5 is 150 (but i'm using my head let me find my calculator)

OpenStudy (brainzonly):

yeah I was right. but why does 150/6 equal 25????

OpenStudy (brainzonly):

OH! I forgot to invert 5/6 LOL

OpenStudy (kevin):

Sorry, I just look on the other site XD

OpenStudy (brainzonly):

Now, (using my head) 30*6 =180...

OpenStudy (brainzonly):

1*5=5, 180/5 equals 30???????????????????????

OpenStudy (brainzonly):

@Kevin Hey! since you abandoned me, I'm abandoning YOU! :(

OpenStudy (kevin):

It's really 25 miles from his start

OpenStudy (brainzonly):

O.O

OpenStudy (kevin):

yup

OpenStudy (brainzonly):

^^I was right the first time. I should start trusting my brain more xD

OpenStudy (kevin):

5/6 x 30 = 25 miles

OpenStudy (brainzonly):

See, I make it too complicated! XD

OpenStudy (kevin):

XD

OpenStudy (brainzonly):

ok, gotta go. A million @brainzonly s to answer! lol

OpenStudy (mathmale):

Kindly focus on helping Livyloo52 to solve this problem, and do not go off on unrelated tangents. Neither of you involved Livyloo52 in any way; you should have.

OpenStudy (mathmale):

@Livyloo52: please involve yourself in this discussion. I imagine you still need more info to be able to solve this problem yourself. Please ask for whatever guidance you need and want.